OpenStudy (elise_a18):

Write the fraction as a percent. Round to the nearest tenth of a percent if necessary. 31/50 0.62% 161.3% 62% 6.2%

OpenStudy (anonymous):

.62

OpenStudy (anonymous):

Take 31 divided by 50

OpenStudy (elise_a18):

yes but i'm not sure between a and c because A is a percent and it has a decimal, but 62 is also a percent without a decimal...

OpenStudy (anonymous):

it would be 62% as 62% = 0.62

OpenStudy (anonymous):

you could type it into a calculator and it will provide you with the answer with the correct decimal

OpenStudy (anonymous):

I can't explain, but that is the answer beautiful.

OpenStudy (anonymous):

or double multiply it be two 62/100 = 62 percent

OpenStudy (elise_a18):

I'm not sure if the answer is suppose to be with the decimal or without.. Without??

OpenStudy (anonymous):

Without a decimal

OpenStudy (anonymous):

if it had the decimal then it'd be less than 1% of 100

OpenStudy (anonymous):

0.62% is sixty two hundredths of a percent, 62% is sixty two percent

OpenStudy (anonymous):

Theyre asking for percent. 31/50=0.62. Move two places to the write and its a percent 62%

OpenStudy (anonymous):

I think everyone else has it, I'm going to go, talk to ya later meh Christian friend @Elise_a18

OpenStudy (elise_a18):

Thanks guys :D i should have told ya'll that i just needed help between 2 answers :3 Talk to ya later Christian Friend Jadey X3

OpenStudy (mathstudent55):

Percent means "divided by 100" or "out of 100." Since you have \(\dfrac{31}{50} \) which has a denominator of 50, it's easy to find the equaivalent fraction with denominator 100. \(\dfrac{31}{50} \times \dfrac{2}{2} = \dfrac{62}{100} \) Now that you have it written as a number divided by 100, that number is the percent. \(\dfrac{31}{50} = 62\%\)